Suppose that pizza delivery times are normally distributed with a population standard deviation of 6 minutes. A random sample of 50 pizza delivery restaurants is taken and has a sample mean delivery time of 36 minutes. Construct a 90% confidence interval for the population mean delivery time.

What is the sample mean?

6

1.87

36

1.64

What is the sample size?

8.33

90

50

36

What is the population standard deviation?

0.85

50

36

6

What is the confidence level?

90%

99%

80%

95%

What is the lower bound of the confidence interval?

34.6

37.4

34.34

34.58

What is the upper bound of the confidence interval?

34.6

37.42

37.66

37.4

What is the error bound margin?

0.85

1.42

1.4

1.64

What is the correct conclusion?

With 90% confidence, we estimate that the true population mean is between 34.58 minutes and 37.42 minutes.

With 90% confidence, we estimate that the true population mean is between 34.34 minutes and 37.66 minutes.

With 90% confidence, we estimate that the true population mean is between 34.6 minutes and 37.4 minutes.
Answer #1

solution Given that ₃ = 36 At n = 50 go% confidence level =1- 90% = 1-0.9 = 0.1 010.05 2 2 24/2 = 20.05 - 1.645 mergin of for- sompie meon is 36 - sample size is 50 - Population standard deviation is 6 - Confidence level is go% - lower bound of the c